An experiment requires 250 mℓ of a 80 μM solution of potassium chloride. A stock solution

of concentration 45 mM is available. What volume of stock solution should be used? Your

answer should be expressed in μℓ to at least 3 significant figures.


1
Expert's answer
2021-10-10T10:25:04-0400

Proportion:

"80 \times 10^{-6} mol" – 1000 mL

x mol – 250 mL

"x = \frac{80 \times 10^{-6} \times 250}{1000} = 20 \times 10^{-6} \;mol" (we need)

Proportion:

"45 \times 10^{-3}" mol – 1000 mL

"20 \times 10^{-6}" mol – y mL

"y = \frac{20 \times 10^{-6} \times 1000}{45 \times 10^{-3}} = 0.444 \;mL"

Answer: 444 μℓ